### v1

# Systematic discrepancies between reference methods for non-covalent interactions within the S66 dataset

Abstract
The accurate treatment of non-covalent interactions is necessary to model a wide range of applications, from molecular crystals to surface catalysts to aqueous solutions and many more. Quantum diffusion Monte Carlo (DMC) and coupled cluster theory with single, double and perturbative triple excitations [CCSD(T)] are considered two widely-trusted methods for treating non-covalent interactions. However, while they have been well-validated for small molecules, recent work has indicated that these two methods can disagree by more than 7.57.57.5\,7.5kcal/mol for larger systems. The origin of this discrepancy remains unknown. Moreover, the lack of systematic comparisons, particularly for medium-sized complexes, has made it difficult to identify which systems may be prone to such disagreements and the potential scale of these differences. In this work, we leverage the latest developments in DMC to compute interaction energies for the entire S66 dataset, containing 66 medium-sized complexes with a balanced representation of dispersion and electrostatic interactions. Comparison to previous CCSD(T) references reveals systematic trends, with DMC predicting stronger binding than CCSD(T) for electrostatic-dominated systems, while the binding becomes weaker for dispersion-dominated systems. We show that the relative strength of this discrepancy is correlated to the ratio of electrostatic and dispersion interactions, as obtained from energy decomposition analysis methods. Finally, we pinpoint systems in the S66 dataset where these discrepancies are particularly prominent, offering cost-effective benchmarks to guide future developments in DMC, CCSD(T) as well as the wider electronic structure theory community.
preprint: AIP/123-QED
I Introduction
Non-covalent interactions play a crucial role in many areas of science. These interactions govern the structure of molecular crystals Beran (2016) (e.g., in pharmaceutical drugs), biomolecules Riley and Hobza (2011) like DNA and proteins and are relevant to supramolecular Raynal et al. (2014) science and nanotechnology. Planas et al. (2013); Poloni, Smit, and Neaton (2012) They also underlie important processes across chemistry and biology, from protein-ligand binding Du et al. (2016), to catalytic reactions, both on the surface R. Rehak et al. (2020) and in solution. Raynal et al. (2014) Understanding and unlocking new processes for these applications will increasingly rely on accurate computational modeling tools that can treat non-covalent interactions. Müller-Dethlefs and Hobza (2000)
Two methods of choice for modeling non-covalent interactions are quantum diffusion Monte Carlo Foulkes et al. (2001) (DMC) and coupled cluster theory Bartlett and Musiał (2007) with single, double and perturbative triple excitations [CCSD(T)]. While these methods may not be as affordable as density functional theory Mattsson et al. (2004) (DFT), the reference data they provide are pivotal for benchmarking and parametrizing the density functional approximations (DFAs) necessary for practical routine simulations. For example, the local density approximation (LDA) and many extensions build upon a DMC-based parametrization of the correlation energy, Ceperley and Alder (1980) while CCSD(T) interaction energy datasets have helped aid in the development of many modern dispersion corrections. Grimme (2006); Grimme et al. (2010, 2016); A. Price, Otero-de-la-Roza, and R. Johnson (2023) In particular, the applicability of these methods to larger systems have rapidly expanded in recent years, arising from computer hardware improvements and, more importantly, algorithmic/methodological developments to both DMC Zen et al. (2016); Krogel (2016); Bennett et al. (2017, 2018); Annaberdiyev et al. (2018); Zen et al. (2019); Needs et al. (2020); Nakano et al. (2020); Kent et al. (2020); Nakano et al. (2023) and CCSD(T). Riplinger and Neese (2013); Riplinger et al. (2016); Ma and Werner (2018); Nagy, Samu, and Kállay (2018); Nagy and Kállay (2019); Masios et al. (2023); Jiang et al. (2024); Ye and Berkelbach (2024)
DMC and CCSD(T) solve the Schrödinger equation to model the systems with distinct approaches and corresponding approximations. Despite these differences, there are many examples where DMC and CCSD(T) have come into alignment. For example, besides small molecules, Dubecký et al. (2013); Řezáč et al. (2015); Raghav et al. (2023) agreement has been obtained for graphene bilayer binding energies Mostaani, Drummond, and Fal’ko (2015), molecular crystal lattice energies, Zen et al. (2018); Della Pia et al. (2022, 2024) molecule-surface interactions Karalti et al. (2012); Al-Hamdani, Alfè, and Michaelides (2017); Tsatsoulis et al. (2017); Al-Hamdani et al. (2017); Brandenburg et al. (2019); Shi et al. (2023) and vacancy formation energies. Shi et al. (2022) Recently, this agreement has been shown to start to falter Al-Hamdani et al. (2021); Schäfer et al. (2024); Ballesteros, Dunivan, and Lao (2021); Villot et al. (2022) for large dispersion-bound molecules, with differences as large as 7.57.57.5\,7.5kcal/mol for a buckyball-ring ($\mathrm{C}_{60}@[6]\mathrm{CPPA}$) complex.
The origin of the discrepancy between DMC and CCSD(T) for large dispersion-bound molecules is a topic of current debate, Schäfer et al. (2024); Fishman et al. (2024); Lambie et al. (2024); Lao (2024) particularly on the validity of the perturbative triples (T) contribution in CCSD(T). Schäfer et al. Schäfer et al. (2024) have suggested that part of this discrepancy arises from missing contributions in (T) that can be accounted by the (cT) approach. In addition, Semidalas et al. Semidalas et al. (2022) have reported non-trivial discrepancies between CCSD(T) and post-CCSD(T) methods such as CCSDT(Q). Conversely, Lambie et al. Lambie et al. (2024) have found that CCSD(T) does not differ significantly against CCSDT(Q) using the Pariser-Parr-Pople (PPP) model Pariser and Parr (1953a, b); Pople (1953) for large conjugated systems. Similarly, Fishman et al. Fishman et al. (2024) and Lao Lao (2024) report only a slight overbinding of CCSD(T) against CCSDT(Q) that cannot explain the discrepancy against DMC.
Understanding these discrepancies between DMC and CCSD(T) for large molecules requires cross-validating these methods across systematic datasets, particularly those involving medium to large sized molecules Sure and Grimme (2015); Ballesteros, Dunivan, and Lao (2021); Grimme (2012) which sample a range of non-covalent interactions. While DMC and CCSD(T) have both been compared (to great agreement) for the A24 Řezáč and Hobza (2013) and S22 Jurečka et al. (2006) datasets of small molecular complexes, DMC has not been frequently applied to study medium-sized datasets. In particular, it has not been used to study the S66 dataset, Řezáč, Riley, and Hobza (2011a) a compilation of 66 dimers that probes the two major types of non-covalent interactions: dispersion and hydrogen-bonding together with those of mixed character. As well as covering a range of interactions, many of the molecules considered form the building blocks for larger biomolecules along different binding configurations. Furthermore, the parallel-displaced benzene dimer Sinnokrot and Sherrill (2004) is included in this set of complexes, making it an interesting modeling challenge. Such a dataset has been pivotal towards benchmarking Goerigk, Kruse, and Grimme (2011); Gao et al. (2016); Peng et al. (2016); Yu and Wang (2020); Grimme et al. (2021); Ehlert et al. (2021); Müller, Hansen, and Grimme (2023); Lu and Chen (2023); Lee et al. (2024) DFAs in DFT as well as lower-level approximations to wave-function methods Řezáč, Riley, and Hobza (2011b); Riley, Řezáč, and Hobza (2012); Altun, Neese, and Bistoni (2020); Shee et al. (2021); Lupi et al. (2021); Semidalas et al. (2022); Beran et al. (2023) and even machine-learning models. Christensen et al. (2021); Villot and Lao (2024)
In this work, we leverage the latest developments in DMC to compute interaction energies for the entire S66 dataset. When compared to CCSD(T) estimates (taken from the literature), we reveal a consistent weaker binding of dispersion interactions and consistent stronger binding of electrostatic interactions in DMC. In particular, we show that their differences are correlated to the ratio of electrostatic and dispersive interactions within the system. The discrepancies in dispersion-dominated systems are shown to be reduced when utilizing an (empirically fitted) CCSD(cT) formulation, Schäfer et al. (2024) although notable differences remain. We identify specific systems with well-defined differences between DMC and CCSD(T) that can serve as model systems for testing future developments in both methods, setting the stage towards resolving their discrepancies.
II Methods
II.1 Diffusion Monte Carlo
The DMC interaction energies of the S66Řezáč, Riley, and Hobza (2011a) dataset are computed as:
$\Delta E_{\textrm{int.}}=E_{\textrm{dimer}}-E_{\textrm{mon. 1}}-E_{\textrm{mon. 2}},$ (1)
where $E_{\mathrm{dimer}}$ is the total energy of the dimer, and $E_{\textrm{mon. 1}},E_{\textrm{mon. 2}}$ are the total energies of the constituent monomers. In the S66 dataset, these monomers are kept fixed to their geometry in the dimer, which is in general different from their equilibrium geometry. In this work, we first computed the energies of the monomers with DMC at a chosen reference geometry. Subsequently, we added the deformation energy, i.e. the energy difference between the geometry of the monomer in the dimer and against this reference geometry using CCSD(T). We provide further details on these calculations in Sec. S2.1 of the supplementary material, and show for a subset of the S66 complexes that differences between DMC and CCSD(T) predictions of the deformation energies are within $\sim 0.12\,$kcal/mol.
A detailed description of the DMC method can be found in Ref. 9. In this work, we compute fixed-node DMC interaction energies by using the CASINO codeNeeds et al. (2020). We use energy-consistent correlated electron pseudopotentialsTrail and Needs (2017) (eCEPP) with the determinant locality approximation (DLA)Zen et al. (2019). The trial wave-functions were of the Slater–Jastrow type with single Slater determinants, and the single-particle orbitals obtained from DFT local-density approximation (LDA) plane-wave calculations performed with PWscfGiannozzi et al. (2009, 2020) using an energy cut-off of 600 Ry and re-expanded in terms of B-splinesAlfè and Gillan (2004). The Jastrow factor included a two-body electron–electron (e–e) term, two-body electron–nucleus (e–n) terms, and three-body electron–electron–nucleus (e–e–n) terms. The variational parameters of the Jastrow have been optimized by minimizing the variance of each system. The final DMC estimates of $\Delta E_{\textrm{int.}}$ were extrapolated towards the zero time step limit ($\tau\to 0$) by making a cubic fit to a series of time step estimates from 0.110.110.11\,0.11au down to 0.0030.0030.003\,0.003au. We have estimated errors on this fit with two approaches: either by taking the $1\sigma$ standard deviation of the fit at zero time step or the difference w.r.t. the zero time step limit given by a linear fit from the two smallest (0.010.010.01\,0.01au and 0.0030.0030.003\,0.003au) time steps; the latter accounts for systems where the time step behavior may change close to the zero time step limit. The final error estimate is taken to be the larger of the two, as discussed in Sec. S2.4 of the supplementary material.
The parameters chosen within the present work follow from previous DMC calculations for large molecules in Ref. 49 as well as molecular crystals in Refs. 40 and 41. Within these studies of non-covalent interactions, the LDA trial wave-function was shown to be valid, either by comparison to experiments or when using trial wave-functions with other DFAs. For the case of the AcOH dimer system (ID 20), we have performed our own validation tests on the choice of trial wave-function as well as localization approximation in Sec. S7 of the supplementary material.
II.2 Coupled Cluster Theory
Several CCSD(T) estimates Řezáč, Riley, and Hobza (2011a, b); Brauer et al. (2016); Kesharwani et al. (2018); Ma and Werner (2019); Altun, Neese, and Bistoni (2020); Lupi et al. (2021); Santra et al. (2022); Semidalas et al. (2022); Nagy et al. (2023) of the S66 interaction energies are available in the literature. Here, we compare DMC to the average of three recent CCSD(T) calculations Řezáč, Riley, and Hobza (2011b); Kesharwani et al. (2018); Nagy et al. (2023): the revised calculations from Řezáč et al.Řezáč, Riley, and Hobza (2011b); the “SILVER” estimates from Kesharwani et al.Kesharwani et al. (2018); and the “14k-GOLD” estimates from Nagy et al.Nagy et al. (2023). A brief description of the three different CCSD(T) calculations is reported in Sec. S3 of the supplementary material. [^1] Note that the difficulty of reaching the complete basis set (CBS) limit is the principle reason for several CCSD(T) evaluations of S66. As reaching the CBS limit remains a challenge, particularly in dispersion bound complexes.
Schäfer et al. Schäfer et al. (2024) have recently demonstrated that there exists an empirical relationship between the (cT) and the (T) correlation contributions to the total energy using the CCSD and MP2 correlation energies. The resulting (cT)-fit is of the form:
$\dfrac{\textrm{(T)}}{\textrm{(cT)-fit}}=a+b\cdot\dfrac{\textrm{MP2 corr.}}{\textrm{CCSD corr.}},$ (2)
where $a$ and $b$ were parameters fitted from comparing CCSD(cT) to CCSD(T) calculations for a set of dispersion-bound complexes. We have recomputed the CCSD(cT)-fit values from the original CCSD(T) “SILVER” estimates from Kesharwani et al., adding the difference between (cT) and (T) to the final (averaged) CCSD(T) estimates, as given in Sec. S4 of the supplementary material.
III Results and Discussion
This dataset comprises of a diverse range of interactions and we have separated the systems according to the original S66 categories of hydrogen-bonded, dispersion-bonded and “mixed”-character systems in Figs. 12 and 3, respectively. The corresponding dimer complexes are visualized in Fig. S1 of the supplementary material. We report the DMC estimate of the interaction energy $\Delta E_{\textrm{int.}}$ above the label of each S66 complex. In all cases, the rrors on $\Delta E_{\textrm{int.}}$ estimates are below 0.260.260.26\,0.26kcal/mol, with the majority below 0.100.100.10\,0.10kcal/mol, facilitating reliable comparisons to other CCSD(T).
Figure 1: Comparison between DMC interaction energies $\Delta E_{\textrm{int.}}$ calculated in the present work against CCSD(T) for a subset of systems in the S66 dataset with hydrogen-bonds. The CCSD(T) estimate is taken as an average from three previous calculations Řezáč, Riley, and Hobza (2011b); Kesharwani et al. (2018); Nagy et al. (2023), with corresponding standard deviation as error. The deviation of CCSD(T) from the DMC is plotted with grey crosses, with the statistical errors (corresponding to one standard deviation $\sigma$). The complex ID and label are provided below the x-axis, while the number above each x-axis tick represents the DMC $\Delta E_{\textrm{int.}}$ estimate, with the error on the last reported digit given in parentheses. The uracil dimer (ID 17) is in its base-pair (BP) configuration.
The strength of $\Delta E_{\textrm{int.}}$ varies significantly across the systems, from as large as $-20.17\pm 0.11\,$kcal/mol for complex 20 (acetic acid dimer) to as weak as $-1.11\pm 0.06\,$kcal/mol for complex 30 (benzene-ethene dimer), being stronger in the H-bonded systems. With gray crosses, we plot the difference between DMC and CCSD(T) estimates (as described in the Methods) for the three classes of interactions. We use DMC as the reference (i.e., zero), and plot blue vertical bars along the horizontal zero-axis representing the errors on the DMC estimates. There is overall excellent agreement, with a mean absolute deviation (MAD) of 0.210.210.21\,0.21kcal/mol across the entire S66 dataset. We find systematic trends in the differences between CCSD(T) and DMC, with CCSD(T) predicting weaker binding compared to DMC for hydrogen-bonded systems in Fig. 1, with an MAD of $\sim 0.24\,$kcal/mol, while predicting a stronger binding for dispersion dominated systems in Fig. 2, with an MAD of $\sim 0.24\,$kcal/mol. For the “mixed” character systems in Fig. 3, the MAD is lower at 0.140.140.14\,0.14kcal/mol.
Figure 2: Comparison between DMC interaction energies $\Delta E_{\textrm{int.}}$ calculated in the present work against CCSD(T) for a subset of systems in the S66 dataset dominated by dispersion interactions. The first 6 dimers are $\pi{-}\pi$ stacked. Refer to the caption of Fig. 1 for the plot details. The CCSD(T) estimate is taken as an average from three previous calculations Řezáč, Riley, and Hobza (2011b); Kesharwani et al. (2018); Nagy et al. (2023), with corresponding standard deviation as error. Additional CCSD(cT)-fit estimates are are reported with golden crosses. These are calculated by scaling the CCSD(T) estimates based on their MP2 and CCSD contributions with the approach described in Ref. 50.
[…]
The stronger binding of DMC over CCSD(T) has not been (systematically) reported before, with the acetic acid dimer (ID 20) giving the maximum deviation of 0.80.80.8\,0.8kcal/mol across all S66 systems. Within Sec. S7 of the supplementary material, we have confirmed that the computed $\Delta E_{\textrm{int.}}$ estimate ($-20.17\pm 0.11\,$kcal/mol) does not depend on the chosen pseudopotential (eCEPP), localization scheme (DLA) or trial wave-function (LDA). For example, we have also performed all-electron calculations, giving an estimate of $-20.19\pm 0.11\,$kcal/mol (at $\tau{=}0.0015\,$au that is within the statistical uncertainties of our original estimate. We have also performed tests using PBE and PBE0 trial wave-functions, showing that the DMC $\Delta E_{\textrm{int.}}$ has a negligible (${<}0.15\,$kcal/mol) dependence on the nodal surface for the DFAs considered. Furthermore, we also computed estimates for two-other localization schemes: T-move and determinant localization T-move, both of which were within the statistical uncertainties of our original estimate. It should be noted that while the absolute value of the difference can be significant for some hydrogen-bonded systems, the relative difference (normalized against the DMC $\Delta E_{\textrm{int.}}$) is significantly smaller, with a mean relative difference of 2.45% compared to 8.17% for the dispersion-dominated systems (see Sec. S6 of the supplementary material).
Figure 4: Error decomposition analysis. We report the absolute difference between DMC and CCSD(T) relative to the DMC values, i.e. $\left(E_{\mathrm{DMC}}-E_{\mathrm{CCSD(T)}}\right)/\left|E_{\mathrm{DMC}}\right|$, as a function of the natural logarithm of the electrostatic (ELST) to dispersion (DISP) ratio contribution to the binding energy. The ELST to DISP ratio is determined from the SAPT analysis from Ref. 93. The color code is red for H-bonded systems (ID from 1 to 23), blue for dispersion dominated systems (ID from 24 to 46), and green for mixed systems (ID from 47 to 66).
The weaker binding of DMC over CCSD(T) for dispersion-dominated systems is now relatively well-documented Al-Hamdani et al. (2021); Schäfer et al. (2024), and there is evidence that it can be improved by replacing the perturbative triples (T) contribution with the recent (cT) Masios et al. (2023) contribution. We plot the difference between an empirically CCSD(cT)-fit formulation as yellow crosses for the dispersion-dominated systems. In all cases, CCSD(cT) has a weaker binding than CCSD(T), leaning closer towards DMC, leading to an MAD of 0.090.090.09\,0.09kcal/mol and mean relative difference of 3.43% However, this does not fully resolve the discrepancies across all of the dispersion systems, with significant discrepancy of ${\sim}0.4\,$kcal/mol remaining for the uracil-cyclopentane dimer (ID 42). Such a significant discrepancy makes this a worthwhile system to investigate further and could give clues on remaining discrepancies between DMC and CCSD(T) observed in Ref. 50. The analysis reported above highlights an important outcome of this work: the identification of smaller, simpler systems that show notable discrepancies between DMC and coupled cluster methods. Specifically, we found a discrepancy of approximately ${\sim}0.9\,$kcal/mol for the hydrogen-bonded acetic acid dimer (ID 20) and ${\sim}0.5\,$kcal/mol for the dispersion-dominated uracil–cyclopentane dimer (ID 43), which contain 64 and 98 (total) electrons, respectively. These medium-sized systems represent an almost tenfold reduction in electron count compared to the larger $\mathrm{C}_{60}@[6]\mathrm{CPPA}$ buckyball-ring system (672 electrons) studied in Ref. 49. Thus, they might offer practical, cost-effective models for further exploring the discrepancy between DMC and coupled-cluster.
Finally, we now focus on the difference between DMC and CCSD(T) as a function of the dispersion and electrostatic contribution to the interaction energy. In particular, we find that the relative differences between DMC and CCSD(T) for each system within the S66 dataset can be correlated to the relative strength of the dispersion and electrostatic interactions that make up its $\Delta E_{\textrm{int.}}$. We used the Symmetry Adapted Perturbation Theory (SAPT) calculations from Burns et al.Burns et al. (2017), which decomposes $\Delta E_{\textrm{int.}}$ into contributions from electrostatics (ELST), exchange, induction, and dispersion (DISP). Notably, we show in Fig. 2 that there is a strong linear trend ($R^{2}{=}0.78$) between the natural logarithm of the ELST and DISP contributions, $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$, and the relative difference (in %) between CCSD(T) and DMC. This analysis confirms our prior observations on the trends between DMC and CCSD(T). For example, the more dominant the DISP contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more negative $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to underbind with respect to DMC. Similarly, the stronger the ELST contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more positive $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to overbind with respect to DMC. We expect that this cheap descriptor can be used in the future to identify more challenging systems with larger discrepancies between DMC and CCSD(T).
IV Conclusions
To summarize, we have computed highly accurate estimates for the S66 dateset – one of the most widely used databases for non-covalent interactions in biological and organic molecules – with fixed-node diffusion quantum Monte Carlo. These estimates have provided new insights into recent discussions on its discrepancies with another widely-trusted method: coupled cluster theory with single, double and perturbative triple excitations [CCSD(T)]. Our data shows systematic trends, with DMC predicting stronger binding in hydrogen-bonded systems than CCSD(T), and weaker binding in dispersion dominated systems. We show that there is a correlation between the relative strength of these discrepancies with the nature of the interaction, specifically the relative ratio of the electrostatic and dispersion contributions to the interaction energy as provided by previous Symmetry Adapted Perturbation Theory (SAPT) calculations. Burns et al. (2017) In addition, we show that the discrepancy between DMC and CCSD(T) on dispersion-dominated systems can be reduced using a recently proposed CCSD(cT) formulation, albeit with still significant remaining differences. While this work does not identify the origin of the disagreement between DMC and CCSD(T), it has identified the type of interactions where it is particularly prevalent and importantly, we have identified model systems within the S66 dataset where these errors are prominent. These results have strong implications for the electronic structure theory community, addressing the knowledge gap on the trends of DMC interaction energies for non-covalent molecular complexes. Furthermore, the accurate reference data produced within this work is expected to benefit the wider materials modeling community, being instrumental for benchmarking applications ranging from the development of machine learned interatomic potentials to crystal structure prediction, drug design, and renewable energy.
Supplementary Material
See the supplementary material for details on the DMC calculations, comprising the convergence of the calculations with respect to the time step, the influence of the choice of the monomer geometry on the dimer interaction energy, as well as tests on the pseudopotential localization error and the Jastrow optimization.
Data Availability
The data that support the findings of this study are available within the article and its supplementary material. All analysis can be found on GitHub (made public upon acceptance of the manuscript) at zenandrea/FNDMC-S66 and can be viewed interactively online through associated Jupyter notebooks (via Google Colab), with links provided in the corresponding GitHub repository.
Acknowledgments
We acknowledge the computational resources from Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service, provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1 and EP/P020259/1), and DiRAC funding from the Science and Technology Facilities Council (www.dirac.ac.uk). We are further grateful for computational support from the UK national high performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and funded by EPSRC grant ref EP/X035891/1. This research also used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. A.M. and B.X.S acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant No. 101071937). D.A. and A.Z. acknowledges support from Leverhulme grant no. RPG-2020-038, and from the European Union under the Next generation EU (projects 20222FXZ33 and P2022MC742).

### v2

@@
 We show that the relative strength of this discrepancy is correlated to the ratio of electrostatic and dispersion interactions, as obtained from energy decomposition analysis methods.
-Finally, we pinpoint systems in the S66 dataset where these discrepancies are particularly prominent, offering cost-effective benchmarks to guide future developments in DMC, CCSD(T) as well as the wider electronic structure theory community.
+Finally, we have pinpointed model systems: the hydrogen-bonded acetic acid dimer (ID 20) and dispersion-dominated uracil–cyclopentane dimer (ID 42), where these discrepancies are particularly prominent.
+These systems offer cost-effective benchmarks to guide future developments in DMC, CCSD(T) as well as the wider electronic structure theory community.
 preprint: AIP/123-QED
@@
 In addition, Semidalas et al.
-Semidalas et al.
-(2022) have reported non-trivial discrepancies between CCSD(T) and post-CCSD(T) methods such as CCSDT(Q).
+Semidalas, Boese, and Martin (2025) have reported non-trivial discrepancies between CCSD(T) and post-CCSD(T) methods such as CCSDT(Q).
 Conversely, Lambie et al.
@@
 The variational parameters of the Jastrow have been optimized by minimizing the variance of each system.
-The final DMC estimates of $\Delta E_{\textrm{int.}}$ were extrapolated towards the zero time step limit ($\tau\to 0$) by making a cubic fit to a series of time step estimates from 0.110.110.11\,0.11au down to 0.0030.0030.003\,0.003au.
-We have estimated errors on this fit with two approaches: either by taking the $1\sigma$ standard deviation of the fit at zero time step or the difference w.r.t.
-the zero time step limit given by a linear fit from the two smallest (0.010.010.01\,0.01au and 0.0030.0030.003\,0.003au) time steps; the latter accounts for systems where the time step behavior may change close to the zero time step limit.
-The final error estimate is taken to be the larger of the two, as discussed in Sec.
+The final DMC estimates of $\Delta E_{\text{int.}}$ were extrapolated towards the zero time step limit ($\tau\to 0$) by making a cubic fit to a series of time step estimates from 0.10.10.1\,0.1au down to 0.0030.0030.003\,0.003au.
+We estimate errors which capture both the stochastic errors in the fit as well as the errors in the cubic fit due to the changing behavior near the zero time step limit.
+To do this, we make a linear fit on a subset of time steps below (and including) 0.020.020.02\,0.02au and calculate the difference of the extrapolated estimates from the linear fit against the original cubic fit.
+The final error estimate is taken to be the larger of the stochastic errors of the cubic fit or the difference between the linear fit and cubic fit, as discussed in Sec.
 S2.4 of the supplementary material.
@@
 Schäfer et al.
-(2024) have recently demonstrated that there exists an empirical relationship between the (cT) and the (T) correlation contributions to the total energy using the CCSD and MP2 correlation energies.
+(2024) have recently demonstrated that there exists an empirical relationship for dispersion-dominated complexes between the (cT) and the (T) correlation contributions to the total energy using the CCSD and MP2 correlation energies.
 The resulting (cT)-fit is of the form:
@@
 III Results and Discussion
+We report the final DMC estimates for the entire S66 dataset in Figs.
+12 and 3.
 This dataset comprises of a diverse range of interactions and we have separated the systems according to the original S66 categories of hydrogen-bonded, dispersion-bonded and “mixed”-character systems in Figs.
@@
 We report the DMC estimate of the interaction energy $\Delta E_{\textrm{int.}}$ above the label of each S66 complex.
-In all cases, the rrors on $\Delta E_{\textrm{int.}}$ estimates are below 0.260.260.26\,0.26kcal/mol, with the majority below 0.100.100.10\,0.10kcal/mol, facilitating reliable comparisons to other CCSD(T).
+In all cases, the errors on $\Delta E_{\textrm{int.}}$ estimates are below 0.120.120.12\,0.12kcal/mol, with the majority below 0.100.100.10\,0.10kcal/mol, facilitating reliable comparisons to CCSD(T).
 Figure 1: Comparison between DMC interaction energies $\Delta E_{\textrm{int.}}$ calculated in the present work against CCSD(T) for a subset of systems in the S66 dataset with hydrogen-bonds.
@@
 The uracil dimer (ID 17) is in its base-pair (BP) configuration.
-The strength of $\Delta E_{\textrm{int.}}$ varies significantly across the systems, from as large as $-20.17\pm 0.11\,$kcal/mol for complex 20 (acetic acid dimer) to as weak as $-1.11\pm 0.06\,$kcal/mol for complex 30 (benzene-ethene dimer), being stronger in the H-bonded systems.
+The strength of $\Delta E_{\textrm{int.}}$ varies significantly across the systems, from as large as $-20.17\pm 0.07\,$kcal/mol for complex 20 (acetic acid dimer) to as weak as $-1.11\pm 0.06\,$kcal/mol for complex 30 (benzene-ethene dimer), being stronger in the H-bonded systems.
 With gray crosses, we plot the difference between DMC and CCSD(T) estimates (as described in the Methods) for the three classes of interactions.
@@
 (2023), with corresponding standard deviation as error.
-Additional CCSD(cT)-fit estimates are are reported with golden crosses.
+Additional CCSD(cT)-fit estimates are reported with golden crosses.
 These are calculated by scaling the CCSD(T) estimates based on their MP2 and CCSD contributions with the approach described in Ref.
@@
 Within Sec.
-S7 of the supplementary material, we have confirmed that the computed $\Delta E_{\textrm{int.}}$ estimate ($-20.17\pm 0.11\,$kcal/mol) does not depend on the chosen pseudopotential (eCEPP), localization scheme (DLA) or trial wave-function (LDA).
-For example, we have also performed all-electron calculations, giving an estimate of $-20.19\pm 0.11\,$kcal/mol (at $\tau{=}0.0015\,$au that is within the statistical uncertainties of our original estimate.
+S7 of the supplementary material, we have confirmed that the computed $\Delta E_{\textrm{int.}}$ estimate ($-20.17\pm 0.07\,$kcal/mol) does not depend on the chosen pseudopotential (eCEPP), localization scheme (DLA) or trial wave-function (LDA).
+For example, we have also performed all-electron calculations, giving an estimate of $-20.32\pm 0.12\,$kcal/mol.
 We have also performed tests using PBE and PBE0 trial wave-functions, showing that the DMC $\Delta E_{\textrm{int.}}$ has a negligible (${<}0.15\,$kcal/mol) dependence on the nodal surface for the DFAs considered.
 Furthermore, we also computed estimates for two-other localization schemes: T-move and determinant localization T-move, both of which were within the statistical uncertainties of our original estimate.
-It should be noted that while the absolute value of the difference can be significant for some hydrogen-bonded systems, the relative difference (normalized against the DMC $\Delta E_{\textrm{int.}}$) is significantly smaller, with a mean relative difference of 2.45% compared to 8.17% for the dispersion-dominated systems (see Sec.
+It should be noted that while the absolute value of the difference can be significant for some hydrogen-bonded systems, the relative difference (normalized against the DMC $\Delta E_{\textrm{int.}}$) is significantly smaller, with a mean relative difference of 2.45% compared to 8.21% for the dispersion-dominated systems (see Sec.
 S6 of the supplementary material).
 Figure 4: Error decomposition analysis.
-We report the absolute difference between DMC and CCSD(T) relative to the DMC values, i.e.
+We report the difference between DMC and CCSD(T) relative to the DMC magnitude, i.e.
 $\left(E_{\mathrm{DMC}}-E_{\mathrm{CCSD(T)}}\right)/\left|E_{\mathrm{DMC}}\right|$, as a function of the natural logarithm of the electrostatic (ELST) to dispersion (DISP) ratio contribution to the binding energy.
 The ELST to DISP ratio is determined from the SAPT analysis from Ref.
-93.
+94.
 The color code is red for H-bonded systems (ID from 1 to 23), blue for dispersion dominated systems (ID from 24 to 46), and green for mixed systems (ID from 47 to 66).
@@
 We plot the difference between an empirically CCSD(cT)-fit formulation as yellow crosses for the dispersion-dominated systems.
-In all cases, CCSD(cT) has a weaker binding than CCSD(T), leaning closer towards DMC, leading to an MAD of 0.090.090.09\,0.09kcal/mol and mean relative difference of 3.43% However, this does not fully resolve the discrepancies across all of the dispersion systems, with significant discrepancy of ${\sim}0.4\,$kcal/mol remaining for the uracil-cyclopentane dimer (ID 42).
+In all cases, CCSD(cT) has a weaker binding than CCSD(T), leaning closer towards DMC, leading to an MAD of 0.090.090.09\,0.09kcal/mol and mean relative difference of 3.44% However, this does not fully resolve the discrepancies across all of the dispersion systems, with significant discrepancy of ${\sim}0.5\,$kcal/mol remaining for the uracil-cyclopentane dimer (ID 42).
 Such a significant discrepancy makes this a worthwhile system to investigate further and could give clues on remaining discrepancies between DMC and CCSD(T) observed in Ref.
@@
 The analysis reported above highlights an important outcome of this work: the identification of smaller, simpler systems that show notable discrepancies between DMC and coupled cluster methods.
-Specifically, we found a discrepancy of approximately ${\sim}0.9\,$kcal/mol for the hydrogen-bonded acetic acid dimer (ID 20) and ${\sim}0.5\,$kcal/mol for the dispersion-dominated uracil–cyclopentane dimer (ID 43), which contain 64 and 98 (total) electrons, respectively.
+Specifically, we found a discrepancy of approximately ${\sim}0.8\,$kcal/mol for the hydrogen-bonded acetic acid dimer (ID 20) and ${\sim}0.5\,$kcal/mol for the dispersion-dominated uracil–cyclopentane dimer (ID 42), which contain 64 and 98 (total) electrons, respectively.
 These medium-sized systems represent an almost tenfold reduction in electron count compared to the larger $\mathrm{C}_{60}@[6]\mathrm{CPPA}$ buckyball-ring system (672 electrons) studied in Ref.
@@
 Thus, they might offer practical, cost-effective models for further exploring the discrepancy between DMC and coupled-cluster.
-Finally, we now focus on the difference between DMC and CCSD(T) as a function of the dispersion and electrostatic contribution to the interaction energy.
+We now focus on the difference between DMC and CCSD(T) as a function of the dispersion and electrostatic contribution to the interaction energy.
 In particular, we find that the relative differences between DMC and CCSD(T) for each system within the S66 dataset can be correlated to the relative strength of the dispersion and electrostatic interactions that make up its $\Delta E_{\textrm{int.}}$.
-We used the Symmetry Adapted Perturbation Theory (SAPT) calculations from Burns et al.Burns et al.
-(2017), which decomposes $\Delta E_{\textrm{int.}}$ into contributions from electrostatics (ELST), exchange, induction, and dispersion (DISP).
+We used the Symmetry Adapted Perturbation Theory (SAPT) calculations (at the $s$SAPT0 level with the jun-cc-pVDZ basis set Parker et al.
+(2014)) from Burns et al.Burns et al.
+(2017), which decomposes $\Delta E_{\text{int.}}$ into contributions from electrostatics (ELST), exchange, induction, and dispersion (DISP).
 Notably, we show in Fig.
-2 that there is a strong linear trend ($R^{2}{=}0.78$) between the natural logarithm of the ELST and DISP contributions, $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$, and the relative difference (in %) between CCSD(T) and DMC.
-This analysis confirms our prior observations on the trends between DMC and CCSD(T).
-For example, the more dominant the DISP contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more negative $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to underbind with respect to DMC.
+4 that there is a strong linear trend ($R^{2}{=}0.78$) between the natural logarithm of the ELST and DISP contributions, $\log\left(\frac{\text{ELST}}{\text{DISP}}\right)$, and the relative difference (in %) between CCSD(T) and DMC.
+In Sec.
+S9 of the supplementary material, we show that this strong linear trend remains at the more sophisticated SAPT2+(3)(CCD)/aug-cc-pVTZ level.
+Villot and Lao (2024) This analysis confirms our prior observations on the trends between DMC and CCSD(T).
+For example, the more dominant the DISP contribution to $\Delta E_{\text{int.}}$ (i.e., a more negative $\log\left(\frac{\text{ELST}}{\text{DISP}}\right)$), the more CCSD(T) is found to underbind with respect to DMC.
 Similarly, the stronger the ELST contribution to $\Delta E_{\textrm{int.}}$ (i.e., a more positive $\log\left(\frac{\textrm{ELST}}{\textrm{DISP}}\right)$), the more CCSD(T) is found to overbind with respect to DMC.
 We expect that this cheap descriptor can be used in the future to identify more challenging systems with larger discrepancies between DMC and CCSD(T).
+Finally, we discuss briefly the potential origins of the observed discrepancies between DMC and CCSD(T) based upon the current literature.
+For H-bonded systems, CCSDT(Q) estimates are available for the A24 dataset Řezáč and Hobza (2013); Řezáč et al.
+(2015) – a set of $\Delta E_{\text{int.}}$ for small dimer complexes.
+Nakano et al.
+Nakano et al.
+(2024) have performed DMC calculations for the entire A24 dataset, where there is a notable discrepancy of $0.26{\pm}0.07\,$kcal/mol and $0.34{\pm}0.07\,$kcal/mol for the water-ammonia and HCN dimer complexes, in line with our observation of stronger binding in DMC.
+The CCSDT(Q) references, albeit at small basis sets, find negligible (${<}0.01\,$kcal/mol) changes relative to CCSD(T).
+For dispersion-bound systems, there exist CCSDT(Q) estimates for the parallel-displaced (PD) benzene dimer by Semidalas et al.
+Semidalas, Boese, and Martin (2025) and by Karton and Martin Karton and Martin (2021), which report differences to CCSD(T) of $-0.085\,$kcal/mol and $-0.058\,$kcal/mol, respectively, using small truncated double-zeta quality basis sets.
+The reported difference between DMC and CCSD(T) is $-0.37{\pm}0.08\,$kcal/mol for this system which indicates that the majority of this difference is not covered when going to CCSDT(Q).
+For both H-bonded and dispersion-bound systems (as well as those of mixed-character), the observed differences could arise either from higher order excitations and larger basis sets needed from coupled cluster theory, or biases in the FN-DMC evaluations, likely coming from the fixed-node approximation.
 IV Conclusions
@@
 The data that support the findings of this study are available within the article and its supplementary material.
-All analysis can be found on GitHub (made public upon acceptance of the manuscript) at zenandrea/FNDMC-S66 and can be viewed interactively online through associated Jupyter notebooks (via Google Colab), with links provided in the corresponding GitHub repository.
+All analysis can be found on GitHub at github.com/zenandrea/FNDMC-S66 and can be viewed interactively online through associated Jupyter notebooks (via Google Colab), with links provided in the corresponding GitHub repository.
 Acknowledgments